OpenStudy (anonymous):

At the Wild Thing Zoo, you can rent a motorized cart to tour the grounds for a $4 initial charge and $4 per hour. An equation that represents the total cost y for x hours is y = 4 + 4x. At Safari Zoo, you can rent the same cart for a $3 inital charge and $5 for hour. An equation that represents the total cost y for x hours is y = 3 + 5x. For how many hours is the total charge the same? please help Answer choices: A. 2 hours B. 3 hours C. 1 hour D. 4 hours

OpenStudy (saifoo.khan):

We have to find after how many hours they will charge the same amount/fee. So for that have to equate both of the equations given.

OpenStudy (saifoo.khan):

y = 4 + 4x AND y = 3 + 5x 4+4x = 3 + 5x Solve for x.

OpenStudy (anonymous):

Well i dont know anything about this im just Helping my kid

OpenStudy (anonymous):

sorry ) : could you still help me

OpenStudy (saifoo.khan):

Sure.Where are you stuck?

OpenStudy (anonymous):

wait i think i got it is one hour

OpenStudy (saifoo.khan):

Yes. You should get 1.

OpenStudy (anonymous):

thank you so much

OpenStudy (saifoo.khan):

You're welcome.

OpenStudy (anonymous):

could you help me with this two question really really fast im in a hurry! 1. x+y=7 4x-y=3 a. (2,5) b. (5,2) c. (2,4) d. (4,2) 2. x+y=3 y=2x-15 a. (5,-5) b. (-10,-5) c. (1,2) d. (6,-3) please help

OpenStudy (saifoo.khan):